solve by substitution
x+2y=11
2x+8y=58

Answer by mananth(16946)

x+2y=11
2x+8y=58

x=11-2y
substitute
2(11-2y)+8y=58

22-4y+8y=58
4y=36
y=9


Plug y in the equation
x+2y=11
x+2*(9)=11
x+18=11
x=-18+11
x=-7